OpenStudy (anonymous):

cubed root of {54 times t^8} cubed root of {54 times t^8} @Mathematics

OpenStudy (anonymous):

\[\sqrt[3]54\times{t ^{8}?}\]

OpenStudy (anonymous):

what is t??

OpenStudy (anonymous):

\[\sqrt[3]{54t^8}=\sqrt[3]{8\times 7\times t^3\times t^3\times t^2}\] \[=2t^2\sqrt[3]{7t^2}\]

OpenStudy (anonymous):

there is no given number for t...54t^11?

OpenStudy (anonymous):

is that cubed?

OpenStudy (anonymous):

@satellite 8*7 does not equals 54

OpenStudy (anonymous):

9x6

OpenStudy (anonymous):

the correct answer is 3*t^2 *{(2t^2)}^1/3

OpenStudy (anonymous):

\[3t ^{2}\sqrt{2t ^{2}1/3}\]

OpenStudy (anonymous):

no no no, the 2t^2 is raised to the power 1/3 not multiplied by it,

OpenStudy (anonymous):

\[3t ^{2}\sqrt{2t ^{1/3}}\] like that?

OpenStudy (anonymous):

no no its cube root of (2t^2 ) multiplied with 3 t^2